# DOS: Directional Object Separation in Text Embeddings for Multi-Object Image Generation

### v3

@@ Abstract
Recent progress in text-to-image (T2I) generative models has led to significant improvements in generating high-quality images aligned with text prompts. However, these models still struggle with prompts involving multiple objects, often resulting in object neglect or object mixing. Through extensive studies, we identify four problematic scenarios, Similar Shapes, Similar Textures, Dissimilar Background Biases, and Many Objects, where inter-object relationships frequently lead to such failures. Motivated by two key observations about CLIP embeddings, we propose DOS (Directional Object Separation), a method that modifies three types of CLIP text embeddings before passing them into text-to-image models. Experimental results show that DOS consistently improves the success rate of multi-object image generation and reduces object mixing. In human evaluations, DOS significantly outperforms four competing methods, receiving 26.24%-43.04% more votes across four benchmarks. These results highlight DOS as a practical and effective solution for improving multi-object image generation. The code will be available at https://github.com/dongnami/DOS.
@@
1 Introduction
Recent progress in text-to-image (T2I) generative models has led to substantial improvements in generating high-quality images that closely align with text prompts (PodellEtAl2023; stabilityai2024stable35; blackforestlabs). These models leverage text embeddings from multiple CLIP (radford2021learning) encoders, and optionally the T5 (raffel2020exploring) encoder, to enhance image-text alignment. However, they still struggle to accurately generate images containing multiple objects. In this work, we identify four scenarios where this issue becomes particularly severe and demonstrate that modifying the CLIP text embeddings, without altering the image generation process within the text-to-image models, can substantially mitigate the problem.
Previous work on multi-object image generation has primarily focused on the relationships between nouns and their associated attributes (rassin2023linguistic; zhuang2024magnet; hu2024token), often overlooking challenges that arise from interactions between multiple noun entities. However, through extensive experiments with state-of-the-art T2I models, we find that failures, specifically object neglect or object mixing, occur frequently in scenarios involving inter-object relationships, even when no attributes are present. To investigate these issues more thoroughly, we focus on prompts that primarily contain objects and avoid descriptive attributes. Following extensive exploration and experimentation, we identify four types of problematic scenarios. As shown in Figure 1, the first three categories represent specific types of inter-object relationships where failures become particularly severe: Similar Shapes, Similar Textures, and Dissimilar Background Biases. Background bias refers to the natural background typically associated with each object (e.g., ‘land’ for ‘deer’ vs. ‘sea’ for ‘sea turtle’). The fourth, Many Objects, captures the compounded effect of numerous inter-object relationships, which may include but are not limited to the first three types.
The motivation for our method stems from two previously reported observations regarding CLIP text embeddings used in T2I models: (1) CLIP employs a causal masking mechanism, where information from earlier tokens is mixed into later token embeddings, leading to information mix-ups in the embeddings of later tokens (zhuang2024magnet; chen2024cat), (2) the difference between two CLIP text embeddings appears to encode directional information, which can be indirectly observed through the behavior of T2I models (hu2024token). Our method aims to mitigate information mix-ups in CLIP text embeddings by constructing directional vectors, referred to as separation vectors, derived from the difference between two CLIP embeddings. These vectors are added to the original text embeddings to inject directional information that encourages the separation of object-specific information.
Across four benchmarks corresponding to the problematic scenarios of Similar Shapes, Similar Textures, Dissimilar Background Biases, and Many Objects, DOS improves the success rate of multi-object image generation while reducing object mixing. In human evaluations, compared to four other models including the baseline, DOS received 26.24%-43.04% more votes than the second-best results across the four benchmarks. Since DOS does not alter the image generation process within the text-to-image model, its inference is approximately four times faster than well-known latent modification methods such as Attend-and-Excite (chefer2023attend). These results demonstrate that DOS is both an effective and efficient algorithm for addressing object neglect and object mixing, which frequently arise from inter-object relationships in multi-object image generation.
2 Related Work
Multi-object image generation involves generating images containing multiple objects, optionally along with their associated attributes. However, during generation, interactions among multiple objects, and optionally their associated attributes, often lead to failures such as object neglect, object mixing, or inaccurate attribute binding. A substantial body of work has been proposed to address these issues (chefer2023attend; hu2024ella; jiang2024comat). Among them, Attend-and-Excite (chefer2023attend) demonstrates that iteratively updating the intermediate image latent can reduce such failures. This inspired a series of follow-up methods based on latent modification, such as CONFORM (meral2024conform). While effective to some extent, these methods typically require expensive iterative gradient updates during inference and often introduce visual artifacts due to deviations from the latent distribution the model was trained on. To overcome these limitations, more recent approaches have focused on modifying only the text embeddings–without altering the intermediate image latent (chen2024cat; hu2024token; zhuang2024magnet). Among these, TEBOpt (chen2024cat) proposes modifying the semantic token embeddings corresponding to each object to better separate them. This approach significantly reduces information bias (i.e., cases where certain objects are more frequently generated than others), but shows only limited improvements in addressing object neglect and object mixing. In contrast, our method leverages directional information to modify all types of CLIP text embeddings–not only semantic token embeddings, but also EOT and pooled embeddings–in a way that promotes object separation, significantly reducing both object neglect and object mixing in multi-object image generation.
@@ 3 Preliminary Analysis
In Figure LABEL:fig:introduction, we identify four types of problematic scenarios involving inter-object relationships: Similar Shapes, Similar Textures, Dissimilar Background Biases, and Many Objects. To examine how each of these scenarios intensifies object neglect or object mixing, we analyze four corresponding aspects in multi-object image generation. For each aspect, we construct two sets of multi-object prompts that contrast in condition. For example, under the shape aspect, the dissimilar shapes condition consists of 20 prompts featuring two objects with dissimilar shapes, following the template “a/an {objectAobject~A} and a/an {objectBobject~B}.” In contrast, the similar shapes condition includes 20 prompts with two similarly shaped objects. The conditions for the other three aspects are defined analogously. Full details on the construction of all prompt sets are provided in Table 6 of Appendix A.
@@ 4 DOS: Directional Object Separation in Text Embeddings
In Figure 2, we observed that certain object pairs are more susceptible to object neglect or object mixing than others. Motivated by this observation, we adjust the strength of the separation vectors for each object pair based on their visual (shape and texture) similarities and background bias dissimilarities. To this end, we first select 42 representative words covering a diverse range of object shapes and textures through interactions with GPT‑o4‑mini‑high (openai_gpt-o4-mini-high_2025) and define prompts of the form $P_{\mathrm{attr},(k,n)}=$ “a $attribute_{k}$} $obj_{n}$}” for all $k\in\{1,2,\dots,42\}$. Similarly, we select 36 representative background phrases and define prompts of the form $P_{\mathrm{bg},(l,n)}=$$background_{l}$}, there is a $obj_{n}$}” for all $l\in\{1,2,\dots,36\}$. We then compare the text embeddings $\bm{c}_{\mathrm{obj/EOT/pool}}^{\mathrm{pure},n}$, obtained from the pure prompt $P_{\mathrm{pure},n}=$ “a $obj_{n}$}”, with embeddings $\bm{c}_{\mathrm{obj/EOT/pool}}^{\mathrm{attr},(k,n)}$ and $\bm{c}_{\mathrm{obj/EOT/pool}}^{\mathrm{bg},(l,n)}$ obtained from the attribute and background prompts, respectively. Cosine similarity is used for this comparison:
@@
for $k\in\{1,2,\dots,42\}$, $l\in\{1,2,\dots,36\}$, and $\tau\in\{\mathrm{obj,EOT,pool}\}$, where $\tau$ denotes the type of text embedding. The resulting similarity vectors $\mathbf{a}_{\tau}^{\mathrm{attr},n}$ and $\mathbf{a}_{\tau}^{\mathrm{bg},n}$ reflect how closely $obj_{n}$ aligns with the predefined visual attributes and background contexts. We compute the same vectors for $obj_{m}$, and define the adaptive strength for the object pair ($obj_{n}$, $obj_{m}$) as:
@@
for $\tau\in\{\mathrm{obj,EOT,pool}\}$, where $\rho(\cdot,\cdot)$ denotes Pearson correlation, and $\sigma(x;x_{\tau,0})$ is a shifted tempered sigmoid function defined as:
@@
Here, the offset $x_{\tau,0}^{\mathrm{attr}}$ is the mean of $\rho(\mathbf{a}_{\tau}^{\mathrm{attr},i},\mathbf{a}_{\tau}^{\mathrm{attr},j})$ over all object pairs $(obj_{i},obj_{j})$ from the MS-COCO (lin2014microsoft) dataset. Similarly, $x_{\tau,0}^{\mathrm{bg}}$ is the mean of $\bigl(1-\rho(\mathbf{a}_{\tau}^{\mathrm{bg},i},\mathbf{a}_{\tau}^{\mathrm{bg},j})\bigr)$ over the same object pairs. These adaptive strengths promote stronger separation for object pairs that are more likely to be mixed or neglected, thereby helping to reduce object neglect and mixing in multi-object image generation. Table 7 in Appendix B.1 provides the full list of 42 representative words describing object shapes and textures and 36 representative background phrases, which are used to construct the prompts $P_{\mathrm{attr},(k,n)}$ and $P_{\mathrm{bg},(l,n)}$, respectively.
@@
Table 1: Quantitative comparison across four benchmarks: Similar Shapes, Similar Textures, Dissimilar Background Biases, and Many Objects. SR denotes Success Rate (↑ higher is better), and MR denotes Mixture Rate (↓ lower is better). The best results are highlighted in bold.
@@
Method
Similar
Shapes
Similar
Textures
Dissimilar
Background Biases
Many
Objects
@@ 5.1 Experimental setup
Our method is based on SDXL (PodellEtAl2023) and has also been implemented on the recently released Stable Diffusion 3.5 (SD3.5) (stabilityai2024stable35). We apply DOS updates to all CLIP text embeddings used by each model. For inference, we use a guidance scale of 5.0 with 50 denoising steps for SDXL, and a guidance scale of 7.0 with 28 steps for SD3.5, which are the official default configurations. The temperature $T$ of the shifted tempered sigmoid functions in Eq. 7 is set to $T=0.6$ for all base models. Additional implementation details are provided in Appendix B.1.
@@
To focus on the four problematic scenarios in multi-object image generation, we construct four corresponding benchmarks: Similar Shapes, Similar Textures, Dissimilar Background Biases, and Many Objects. Each of the first three benchmarks consists of 50 two-object prompts, following the template: “a/an $\{object~A\}$ and a/an $\{object~B\}$.” The object pairs for these benchmarks are selected through interactions with GPT-o4-mini-high (openai_gpt-o4-mini-high_2025). For the Many Objects benchmark, we extend the animal list used in Attend-and-Excite to include 17 animals, from which we randomly sample to create 25 three-object prompts and 25 four-object prompts. For each method, we generate 4 images per prompt using different random seeds, resulting in 200 generated images per benchmark. Detailed information about the construction of each benchmark is provided in Appendix B.2.
@@
We adjust the VLM-based metric from EnMMDiT (wei2024enhancing) to measure both the success rate (SR) and mixture rate (MR) for each generated image. SR evaluates whether all the objects specified in the prompt are clearly present in the image, while MR assesses whether any of the objects appear mixed. Specifically, we provided GPT-4o-mini with the generated image and the text prompt used for image generation, asking it to classify each object in the prompt as (1) fully intact, (2) mixed, or (3) absent. SR is calculated as the ratio of images where all objects are classified as fully intact, and MR is calculated as the ratio of images where one or more objects are classified as mixed. A high SR score and a low MR score indicate superior image-text alignment. Additional details about the VLM-based metrics are provided in Appendix B.3.
@@ 5.2 Experimental Results
We compare our method with TEBOpt (chen2024cat), Attend-and-Excite (A&E) (chefer2023attend), and CONFORM (meral2024conform) using SDXL as the base model. For SD3.5, we only compare our method with the baseline model, as it is newly released and methods based on this architecture are not yet available.
@@ 5.3 Analysis
Our method applies DOS updates to three types of CLIP text embeddings: semantic token embeddings corresponding to object nouns, EOT embeddings, and pooled embeddings. In Table 4, we ablate the type of embeddings to which DOS is applied. The results show that applying DOS to either the semantic token embeddings or the EOT/pooled embeddings leads to notable improvements in SR, while MR is more significantly reduced when DOS is applied to EOT/pooled embeddings. This is likely because EOT and pooled embeddings tend to capture the overall semantics of the prompt (zhuang2024magnet; wu2024relation), influencing the overall spatial structure of the generated images when modified. Finally, applying DOS to all three types of embeddings results in the best performance on both SR and MR. Qualitative examples for this ablation study is shown in Figure 10 of Appendix C.2.
@@
To examine the effect of the adaptive strengths defined in Eq. 5, we perform an ablation by setting all adaptive strengths to a fixed value of 0.5. The results, shown in Table 5, indicate that using the fixed strength for all separation vectors already leads to substantial performance improvements compared to the baseline. This suggests that the modification directions specified by the separation vectors are effective on its own, even without adaptive strength modulation. However, applying adaptive strength results in even better performance, indicating that modulating the strength of separation differently for each object pair further enhances image generation with multiple objects.
Method
Similar
Shapes
Similar
Textures
Dissimilar
Background Biases
Many
Objects
@@
Method
Similar
Shapes
Similar
Textures
Dissimilar
Background Biases
Many
Objects
@@
Figure 6: Example where adaptive strengths fail to properly scale separation vectors, leading to object neglect. Optimal strengths are determined through grid search.

### v4

@@ Abstract
Recent progress in text-to-image (T2I) generative models has led to significant improvements in generating high-quality images aligned with text prompts. However, these models still struggle with prompts involving multiple objects, often resulting in object neglect or object mixing. Through extensive studies, we identify four problematic scenarios, Similar Shapes, Similar Textures, Dissimilar Background Biases, and Many Objects, where inter-object relationships frequently lead to such failures. Motivated by two key observations about CLIP embeddings, we propose DOS (Directional Object Separation), a method that modifies three types of CLIP text embeddings before passing them into text-to-image models. Experimental results show that DOS consistently improves the success rate of multi-object image generation and reduces object mixing. In human evaluations, DOS significantly outperforms four competing methods, receiving 26.24%-43.04% more votes across four benchmarks. These results highlight DOS as a practical and effective solution for improving multi-object image generation.
Figure 1: Text-to-image generative models often struggle to generate images with multiple objects, especially in four scenarios where relationships between objects, referred to as inter-object relationships, frequently lead to object neglect or object mixing: Similar Shapes, Similar Textures, Dissimilar Background Biases, and Many Objects. Our method mitigates these failures by modifying text embeddings before feeding them into the text-to-image model.
Codehttps://github.com/dongnami/DOS
1 Introduction
Recent progress in text-to-image (T2I) generative models has led to substantial improvements in generating high-quality images that closely align with text prompts (Podell et al. 2023; Stability AI 2024; Black Forest Labs 2025). These models leverage text embeddings from multiple CLIP (Radford et al. 2021) encoders, and optionally the T5 (Raffel et al. 2020) encoder, to enhance image-text alignment. However, they still struggle to accurately generate images containing multiple objects. In this work, we identify four scenarios where this issue becomes particularly severe and demonstrate that modifying the CLIP text embeddings, without altering the image generation process within the text-to-image models, can substantially mitigate the problem.
Previous work on multi-object image generation has primarily focused on the relationships between nouns and their associated attributes (Rassin et al. 2023; Zhuang et al. 2024; Hu et al. 2024a), often overlooking challenges that arise from interactions between multiple noun entities. However, through extensive experiments with state-of-the-art T2I models, we find that failures, specifically object neglect or object mixing, occur frequently in scenarios involving inter-object relationships, even when no attributes are present. To investigate these issues more thoroughly, we focus on prompts that primarily contain objects and avoid descriptive attributes. Following extensive exploration and experimentation, we identify four types of problematic scenarios. As shown in Figure 1, the first three categories represent specific types of inter-object relationships where failures become particularly severe: Similar Shapes, Similar Textures, and Dissimilar Background Biases. Background bias refers to the natural background typically associated with each object (e.g., ‘land’ for ‘deer’ vs. ‘sea’ for ‘sea turtle’). The fourth, Many Objects, captures the compounded effect of numerous inter-object relationships, which may include but are not limited to the first three types.
The motivation for our method stems from two previously reported observations regarding CLIP text embeddings used in T2I models: (1) CLIP employs a causal masking mechanism, where information from earlier tokens is mixed into later token embeddings, leading to information mix-ups in the embeddings of later tokens (Zhuang et al. 2024; Chen et al. 2024), (2) the difference between two CLIP text embeddings appears to encode directional information, which can be indirectly observed through the behavior of T2I models (Hu et al. 2024a). Our method aims to mitigate information mix-ups in CLIP text embeddings by constructing directional vectors, referred to as separation vectors, derived from the difference between two CLIP embeddings. These vectors are added to the original text embeddings to inject directional information that encourages the separation of object-specific information.
Across four benchmarks corresponding to the problematic scenarios of Similar Shapes, Similar Textures, Dissimilar Background Biases, and Many Objects, DOS improves the success rate of multi-object image generation while reducing object mixing. In human evaluations, compared to four other models including the baseline, DOS received 26.24%-43.04% more votes than the second-best results across the four benchmarks. Since DOS does not alter the image generation process within the text-to-image model, its inference is approximately four times faster than well-known latent modification methods such as Attend-and-Excite (Chefer et al. 2023). These results demonstrate that DOS is both an effective and efficient algorithm for addressing object neglect and object mixing, which frequently arise from inter-object relationships in multi-object image generation.
2 Related Work
Multi-object image generation involves generating images containing multiple objects, optionally along with their associated attributes. However, during generation, interactions among multiple objects, and optionally their associated attributes, often lead to failures such as object neglect, object mixing, or inaccurate attribute binding. A substantial body of work has been proposed to address these issues (Chefer et al. 2023; Hu et al. 2024b; Jiang et al. 2024). Among them, Attend-and-Excite (Chefer et al. 2023) demonstrates that iteratively updating the intermediate image latent can reduce such failures. This inspired a series of follow-up methods based on latent modification, such as CONFORM (Meral et al. 2024). While effective to some extent, these methods typically require expensive iterative gradient updates during inference and often introduce visual artifacts due to deviations from the latent distribution the model was trained on. To overcome these limitations, more recent approaches have focused on modifying only the text embeddings–without altering the intermediate image latent (Chen et al. 2024; Hu et al. 2024a; Zhuang et al. 2024). Among these, TEBOpt (Chen et al. 2024) proposes modifying the semantic token embeddings corresponding to each object to better separate them. This approach significantly reduces information bias (i.e., cases where certain objects are more frequently generated than others), but shows only limited improvements in addressing object neglect and object mixing. In contrast, our method leverages directional information to modify all types of CLIP text embeddings–not only semantic token embeddings, but also EOT and pooled embeddings–in a way that promotes object separation, significantly reducing both object neglect and object mixing in multi-object image generation.
@@ 3 Preliminary Analysis
In Figure 1, we identify four types of problematic scenarios involving inter-object relationships: Similar Shapes, Similar Textures, Dissimilar Background Biases, and Many Objects. To examine how each of these scenarios intensifies object neglect or object mixing, we analyze four corresponding aspects in multi-object image generation. For each aspect, we construct two sets of multi-object prompts that contrast in condition. For example, under the shape aspect, the dissimilar shapes condition consists of 20 prompts featuring two objects with dissimilar shapes, following the template “a/an {objectAobject~A} and a/an {objectBobject~B}.” In contrast, the similar shapes condition includes 20 prompts with two similarly shaped objects. The conditions for the other three aspects are defined analogously. Full details on the construction of all prompt sets are provided in Table 6 of Appendix A.
@@ 4 DOS: Directional Object Separation in Text Embeddings
In Figure 2, we observed that certain object pairs are more susceptible to object neglect or object mixing than others. Motivated by this observation, we adjust the strength of the separation vectors for each object pair based on their visual (shape and texture) similarities and background bias dissimilarities. To this end, we first select 42 representative words covering a diverse range of object shapes and textures through interactions with GPT‑o4‑mini‑high (OpenAI 2025) and define prompts of the form $P_{\mathrm{attr},(k,n)}=$ “a $attribute_{k}$} $obj_{n}$}” for all $k\in\{1,2,\dots,42\}$. Similarly, we select 36 representative background phrases and define prompts of the form $P_{\mathrm{bg},(l,n)}=$$background_{l}$}, there is a $obj_{n}$}” for all $l\in\{1,2,\dots,36\}$. We then compare the text embeddings $\bm{c}_{\mathrm{obj/EOT/pool}}^{\mathrm{pure},n}$, obtained from the pure prompt $P_{\mathrm{pure},n}=$ “a $obj_{n}$}”, with embeddings $\bm{c}_{\mathrm{obj/EOT/pool}}^{\mathrm{attr},(k,n)}$ and $\bm{c}_{\mathrm{obj/EOT/pool}}^{\mathrm{bg},(l,n)}$ obtained from the attribute and background prompts, respectively. Cosine similarity is used for this comparison:
@@
for $k\in\{1,2,\dots,42\}$, $l\in\{1,2,\dots,36\}$, and $\tau\in\{\mathrm{obj,EOT,pool}\}$, where $\tau$ denotes the type of text embedding. The resulting lists of cosine similarities $\mathbf{a}_{\tau}^{\mathrm{attr},n}$ and $\mathbf{a}_{\tau}^{\mathrm{bg},n}$ reflect how closely $obj_{n}$ aligns with the predefined visual attributes and background contexts. We compute the same lists for $obj_{m}$, and define the adaptive strength for the object pair ($obj_{n}$, $obj_{m}$) as:
@@
for $\tau\in\{\mathrm{obj,EOT,pool}\}$, where $\rho(\cdot,\cdot)$ denotes Pearson correlation, which is used to avoid mean and scaling effects, and $\sigma(x;x_{\tau,0})$ is a shifted tempered sigmoid function defined as:
@@
Here, the offset $x_{\tau,0}^{\mathrm{attr}}$ is the mean of $\rho(\mathbf{a}_{\tau}^{\mathrm{attr},i},\mathbf{a}_{\tau}^{\mathrm{attr},j})$ over all object pairs $(obj_{i},obj_{j})$ from the MS-COCO (Lin et al. 2014) dataset. Similarly, $x_{\tau,0}^{\mathrm{bg}}$ is the mean of $\bigl(1-\rho(\mathbf{a}_{\tau}^{\mathrm{bg},i},\mathbf{a}_{\tau}^{\mathrm{bg},j})\bigr)$ over the same object pairs. These adaptive strengths promote stronger separation for object pairs that are more likely to be mixed or neglected, thereby helping to reduce object neglect and mixing in multi-object image generation. In Appendix B.1, we provide the full list of 42 representative words describing object shapes and textures and 36 representative background phrases, which are used to construct the prompts $P_{\mathrm{attr},(k,n)}$ and $P_{\mathrm{bg},(l,n)}$, respectively.
@@
Table 1: Quantitative comparison across four benchmarks: Similar Shapes, Similar Textures, Dissimilar Background Biases, and Many Objects. SR denotes Success Rate (↑ higher is better), and MR denotes Mixture Rate (↓ lower is better).
Method Similar Shapes Similar Textures Dissimilar Background Biases Many Objects
@@ 5.1 Experimental Setup
Our method is based on SDXL (Podell et al. 2023) and has also been implemented on the recently released Stable Diffusion 3.5 (SD3.5) (Stability AI 2024). We apply DOS updates to all CLIP text embeddings used by each model. For inference, we use a guidance scale of 5.0 with 50 denoising steps for SDXL, and a guidance scale of 7.0 with 28 steps for SD3.5, which are the official default configurations. The temperature $T$ of the shifted tempered sigmoid functions in Eq. 7 is set to $T=0.6$ for all base models, determined through a light validation process. Additional implementation details are provided in Appendix B.1.
@@
To focus on the four problematic scenarios in multi-object image generation, we construct four corresponding benchmarks: Similar Shapes, Similar Textures, Dissimilar Background Biases, and Many Objects. Each of the first three benchmarks consists of 50 two-object prompts, following the template: “a/an $\{object~A\}$ and a/an $\{object~B\}$.” The object pairs for these benchmarks are selected through interactions with GPT-o4-mini-high (OpenAI 2025). For the Many Objects benchmark, we extend the animal list used in Attend-and-Excite to include 17 animals, from which we randomly sample to create 25 three-object prompts and 25 four-object prompts. For each method, we generate 4 images per prompt using different random seeds, resulting in 200 generated images per benchmark. While the main experiment for Many Objects focuses on the animal list, we present results for an extended benchmark in Table 11 of Appendix C.3 that also includes electronics, vehicles, and fruits. Detailed information about the construction of each benchmark is provided in Appendix B.2.
@@
We adjust the VLM-based metric from EnMMDiT (Wei et al. 2024) to measure both the success rate (SR) and mixture rate (MR) for each generated image. SR evaluates whether all the objects specified in the prompt are clearly present in the image, while MR assesses whether any of the objects appear mixed. Specifically, we provided GPT-4o-mini with the generated image and the text prompt used for image generation, asking it to classify each object in the prompt as (1) fully intact, (2) mixed, or (3) absent. SR is calculated as the ratio of images where all objects are classified as fully intact, and MR is calculated as the ratio of images where one or more objects are classified as mixed. A high SR score and a low MR score indicate superior image-text alignment. We further validate the consistency of the results across two additional evaluators in Table 13 of Appendix C.6. Additional details about the VLM-based metrics are provided in Appendix B.3.
@@ 5.2 Experimental Results
We compare our method with TEBOpt (Chen et al. 2024), Attend-and-Excite (A&E) (Chefer et al. 2023), and CONFORM (Meral et al. 2024) using SDXL as the base model. For SD3.5, we only compare our method with the baseline model in our main experiment. Further comparisons on SD3.5 against recent methods, TEBOpt (Chen et al. 2024) and Self-Cross (Qiu et al. 2025), are provided in Table 14 of Appendix C.7.
@@ 5.3 Analysis
Our method applies DOS updates to three types of CLIP text embeddings: semantic token embeddings corresponding to object nouns, EOT embeddings, and pooled embeddings. In Table 4, we ablate the type of embeddings to which DOS is applied. The results show that applying DOS to either the semantic token embeddings or the EOT/pooled embeddings leads to notable improvements in SR, while MR is more significantly reduced when DOS is applied to EOT/pooled embeddings. This is likely because EOT and pooled embeddings tend to capture the overall semantics of the prompt (Wu et al. 2024), influencing the overall spatial structure of the generated images when modified. Finally, applying DOS to all three types of embeddings results in the best performance on both SR and MR. Qualitative examples for this ablation study are shown in Figure 10 of Appendix C.2.
@@
To examine the effect of the adaptive strengths defined in Eq. 5, we perform an ablation by setting all adaptive strengths to a fixed value of 0.5. The results, shown in Table 5, indicate that using the fixed strength for all separation vectors already leads to substantial performance improvements compared to the baseline. This suggests that the modification directions specified by the separation vectors are effective on its own, even without adaptive strength modulation. However, applying adaptive strength results in even better performance, indicating that modulating the strength of separation differently for each object pair further enhances image generation with multiple objects. An additional ablation study is provided in Table 12 of Appendix C.4, examining sensitivity to predefined words and phrases used to compute adaptive strengths.
Method Similar Shapes Similar Textures Dissimilar Background Biases Many Objects
@@
Method Similar Shapes Similar Textures Dissimilar Background Biases Many Objects
@@
Figure 6: Example where adaptive strengths fail to properly scale separation vectors, leading to object neglect.